What is the df and critical value of F for each situation below?

    a) K=6, n=7 (between-subject design)            _______________________________

               

                b) K=4, n=11(between-subject design)           _______________________________

               

                c) K=3, n=5(between-subject design)             _______________________________
Answer #1

here level of significance not mentioned , considering it 0.05

a(

degree of freedom of numerator =k-1=5

degree of freedom of Denominator =nk-k=7*6*6 =36

critical value of F =2.477

b)

degree of freedom of numerator =k-1=3

degree of freedom of Denominator =nk-k=4*11-4=40

critical value of F =2.839

c)

degree of freedom of numerator =k-1=2

degree of freedom of Denominator =nk-k=3*5-3 =12

critical value of F =3.885